Can you help me fill the empty spots in this table? Please show me how to do it. Remember that concentrate hydrochloric acid is a 12.1 M solution of HCl in waterCompound Structure MW (g/mol) Mass (g) mols Density g/mL 1.000 @ 3.9 deg Volume (mL) .3 mL 18.02 H2O Not required Water CAS #
Answer #1
Compound Density (g/mL ) MW (g/mol) Mass (g) Volume (mL)   Moles
Water 1.0 g/mL 18.02

300 mg

/0.300 g

0.3 mL

Not required

Maleic acid

1.59 g/mL

116.08

200 mg/

(0.200 g)

0.125 mL

0.00172 mol

Fumaric acid

Not required

116.08

Not required

Not required

Not required

Hydrochloric acid

Not required

36.461 0.110 g 0.25 mL 0.003025 mol

HCl

Molarity =12.1 M

Ethylacetate

Not required

88.1

Not required

Not required

Not required

Absolute Ethanol

Not required

46.08

Not required

Not required

Not required

Let's see bold figure, how they are filled..

1) We know that ,

Mass = volume x density

So by putting values, mass of water = 0.3 mL x 1.0 g/mL

= 0.300 g of water

2)

maleic acid moles and volume has been calculated as..

We know that ,

mass of Maleic acid = 200 mg

We can convert mg into gram by dividing with 1000 , so

= (200/1000)

= 0.200g of Maleic acid

# density of Maleic acid =1.59 g/mL

# mass of Maleic acid =0.200 g

We know that,

(Mass /density ) = volume

So,

(0.200 g / 1.59 g/mL ) = 0.125 mL

# moles of Maleic acid = (mass of Maleic acid / MW of Maleic acid)

= (0.200 g / 116.08 g/mol)

= 0.00172 mol

3) HCl solution -

#Molarity of HCl =12.1 M

#Volume of HCl solution =0.25 mL

We have to convert mL into liter by dividing 1000 ,

So , (0.25 /1000) = 0.00025 liter of HCl

Now we know definition of

Molarity = number of moles (n) /per liter of solution

So by rearranging,

Number of moles (n) = Molarity (M) x volume of solution (L)

=12.1 x 0.00025

=0.003025 mol